मराठी

The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved

Advertisements
Advertisements

प्रश्न

The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved, if they try independently?

बेरीज
Advertisements

उत्तर

Let A be the event that X solves the problem B be the event that Y solves the problem.
Since the odds against student X solving the problem are 8: 6
∴ Probability of occurrence of event A is given by.

P(A) = `6/(8 + 6) = 6/14` and

P(A') = 1 – P(A) = `1 - 6/14 = 8/14`

Also, the odds in favour of student Y solving the problem are 14: 16
∴ Probability of occurrence of event B is given by

P(B) = `14/(14 + 16) = 14/30` and

P(B') = 1 – P(B) = `1 - 14/30 = 16/30`

Now A and B are independent events.
∴ A' and B' are independent events.
A' ∩ B' = Event that neither solves the problem
= P(A' ∩ B') = P(A') · P(B')

= `8/14 xx 16/30`

= `32/105`
A ∪ B = the event that the problem is solved
∴ P (problem will be solved) = P(A ∪ B)
= 1 – P(A ∪ B)`
= 1 – P(A' ∩ B')

= `1 - 32/105`

= `73/105`

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 7: Probability - Miscellaneous Exercise 7 [पृष्ठ ११०]

APPEARS IN

बालभारती Mathematics and Statistics (Commerce) Part 2 [English] Standard 11 Maharashtra State Board
पाठ 7 Probability
Miscellaneous Exercise 7 | Q 14. (a) | पृष्ठ ११०

संबंधित प्रश्‍न

A bag contains 4 balls. Two balls are drawn at random (without replacement) and are found to be white. What is the probability that all balls in the bag are white?


If P(A) = 0·4, P(B) = p, P(A ⋃ B) = 0·6 and A and B are given to be independent events, find the value of 'p'.


A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, problem is solved?


The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the probability that neither solves the problem?


A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2` and `5/8`. Find the probability that the target

  1. is hit exactly by one of them
  2. is not hit by any one of them
  3. is hit
  4. is exactly hit by two of them

A bag contains 3 red and 5 white balls. Two balls are drawn at random one after the other without replacement. Find the probability that both the balls are white.

Solution: Let,

A : First ball drawn is white

B : second ball drawn in white.

P(A) = `square/square`

After drawing the first ball, without replacing it into the bag a second ball is drawn from the remaining `square` balls.

∴ P(B/A) = `square/square`

∴ P(Both balls are white) = P(A ∩ B)

`= "P"(square) * "P"(square)`

`= square * square`

= `square`


Solve the following:

If P(A) = `"P"("A"/"B") = 1/5, "P"("B"/"A") = 1/3` the find `"P"("B'"/"A'")`


Two dice are thrown together. Let A be the event ‘getting 6 on the first die’ and B be the event ‘getting 2 on the second die’. Are the events A and B independent?


If A and B′ are independent events then P(A′ ∪ B) = 1 – ______.


Three events A, B and C have probabilities `2/5, 1/3` and `1/2`, , respectively. Given that P(A ∩ C) = `1/5` and P(B ∩ C) = `1/4`, find the values of P(C|B) and P(A' ∩ C').


Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: 1 – (1 – P1)(1 – P2) 


If A and B are two events such that P(B) = `3/5`, P(A|B) = `1/2` and P(A ∪ B) = `4/5`, then P(A) equals ______.


Two events 'A' and 'B' are said to be independent if


If P(A) = `3/5` and P(B) = `1/5`, find P(A ∩ B), If A and B are independent events.


A problem in Mathematics is given to three students whose chances of solving it are `1/2, 1/3, 1/4` respectively. If the events of their solving the problem are independent then the probability that the problem will be solved, is ______.


What are independent events in probability?


Two events E and F are independent when which conditional probability condition holds, provided \[P(E) \ne 0\]?


For \[E=\{3,6\}\] and \[F=\{2,4,6\}\], what is \[E\cap F\]?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×